Chlorothiazide can be used to treat moderate cases of hypertension. An average oral dose consists of 500 mg. Suppose a patient is given chlorothiazide suspension that has a concentration of 50 mg/ml. How many complete doses are contained in a 237 ml bottle?
Answer #1

Ordered dose=500 mg

1 ml suspension contains 50 mg

So, 237 ml= 237× 50 =11850 mg

Therefore, complete doses in 237 ml bottle= total dose / ordered dose=11850÷500=23.7 = 23 Complete dose